If Maria hit 8 home runs in 15 softball games, then she should hit 24
home runs in 45.
help again

Answer by checkley71(8403)

8/15=.53333 HOMERS PER GAME IS HER RATE.
24/45=.53333 HOMERS PER GAME @ THIS RATE ALSO.
3 TIMES AS MANY HOMERS IN 3 TIMES AS MANY GAMES THUS THE RATE IS THE SAME.
(3*8)=24
(3*15)=45